# Complementarity-based complementarity

### v2

@@ Abstract
We show that the amount of complementarity that a quantum system can exhibit depends on which complementary properties one is considering. Consider a set of mutually unbiased bases (MUBs) which corresponds to the spectral decomposition of some maximally complementary properties: The perfect knowledge of one, i.e. the state of the system is one of the basis states, implies that all the others are completely unknown, namely the measurement of another property will find any of its possible outcomes with uniform probability. We show that, depending on which three of the MUBs we choose, a 5-dimensional system can have different degrees of complementarity, measured using the minimum over all the system states of the sum of the entropies and of the variances of the measurement outcomes related to the three chosen observables. This property was first found experimentally, and demonstrates that an experimental discovery can disclose quantum information effects.
@@
Up to now a symmetry among different complementary properties (MUBs) was implicitly assumed. It was thought that a permutation of the MUBs would not change the information that one could have on them. In this paper we show that this belief is mistaken, by providing a counterexample in dimension $d=5$ (where 6 MUBs exist). The information that a system can possess on three of the six complementary properties depends on which of the properties (MUBs) are considered. We show this by calculating the lower bound of the sum of the entropies or of the variances of the measurement outcomes of three complementary observables, and prove that such a lower bound depends explicitly on the choice of complementary observables: Complementarity-based complementarity (Fig. 1). The analytic results are complemented by an experimental verification and by Monte-Carlo simulations where we calculate the sum of entropies and variances of random states.
The effect reported in this paper was first discovered experimentally while testing entropic uncertainty relations (EURs) in large dimensional Hilbert spaces.
@@
Figure 1: Alice and Bob are both interested in joint values of observables with mutually unbiased bases (MUB) as eigenstates. In dimension 5 there are 6 such observables: Bob measures observables $A,B,C$, Alice measures $D,E,F$. Surprisingly, Alice gets more information: she can divide the $DEF$ phase space in smaller uncertainty blocks than Bob’s, even though they are both looking at MUBs: the amount of complementarity among observables depend on which complementary observables, $ABC$ or $DEF$, are considered.
I Complementarity of three MUBs
@@
Consider three maximally complementary observables $\hat{A},\hat{B},\hat{C}$ with mutually-unbiased eigenstates $|a_{i}\rangle,|b_{j}\rangle,|c_{k}\rangle$. One can have partial knowledge of all three of them if the state $|\psi\rangle$ is a nontrivial superposition when expressed in any of the bases $|a_{i}\rangle,|b_{j}\rangle,|c_{k}\rangle$. How much joint information on them can one obtain? We need a quantification of how uncertain the outcomes of measurements of all three observables are. In the following, we concentrate on two quantifiers  [^1] We also considered the Renyi entropies of order 2 and 1/2, but the numerics gave inconclusive results. On the one hand, we consider the sum of the Shannon entropies $H(\hat{O})=-\sum_{j}p_{j}\log_{2}p_{j}$ of the Born probabilities $p_{j}=|\langle o_{j}|\psi\rangle|^{2}$ pertaining to the measurement of $\hat{O}$. The entropies are all positive quantities, so a small sum implies large joint knowledge. Moreover, the sum of entropies for maximally complementary observables has always a nontrivial lower bound given by the entropic uncertainty relations (EUR) [7, 8, 9]. On the other hand, we consider the sum of the variances [10] of $\hat{A},\hat{B},\hat{C}$. Differently from the entropy, the variance depends also on the eigenvalues of the observables, not only on the probabilities. We will choose observables with eigenvalues equal to a permutation $P(i)$ of the basis index $i$, i.e. $\hat{A}=\sum_{i}P(i)|a_{i}\rangle\langle a_{i}|$, and then minimize the variance $\Delta A^{2}=\langle\hat{A}^{2}\rangle-\langle\hat{A}\rangle^{2}$ over the permutations, to avoid effects due to the arbitrariness of the eigenvalue assignments. Also the sum of variances has a nonzero  [^2] Instead, the product of variances, as in the traditional Heisenberg-Robertson uncertainty relations [24], is problematic because it has no nonzero lower bound when minimizing over states: the product is null already if only one of the two variances is nonzero. lower bound for maximally complementary observables [12, 13].
We now show that the lower bound on the sum of entropies or variances of three different maximally complementary observables $\hat{A},\hat{B},\hat{C}$ depends on the choice of the maximally complementary observables, overturning an implicit belief that all MUBs are equivalent. This uncovers a hitherto unknown asymmetry in high dimensional systems.
@@
We start from dimension $d=5$ where there exist six maximally complementary observables $\hat{A},\hat{B},\hat{C},\hat{D},\hat{E},\hat{F}$ with eigenvectors equal to the six MUBs [14, 15] $|a_{i}\rangle,|b_{j}\rangle,|c_{k}\rangle,|d_{l}\rangle,|e_{m}\rangle,|f_{n}\rangle$ presented in the appendix. In this case, we find that
@@
The bound (1) was known [9] and was implicitly assumed to hold for every possible choice of MUB triplet, whereas the bound (2), obtained numerically (see in the appendix), was found while experimentally checking the entropic uncertainty relations. The states that minimize the first bound are eigenstates of any of the three $\hat{A},\hat{B},\hat{C}$, e.g. $|\psi\rangle=|a_{0}\rangle$. The states that minimize the second bound, instead, have a single null component when expressed in the computational basis $|a_{i}\rangle$, e.g. a state $\sum_{n}\psi_{n}e^{i\phi_{n}}|a_{n}\rangle$ with $\{\psi_{n}\}=(0.54488,0.45067,0,0.45067,0.54488)$, $\{\phi_{n}\}=\left(0,-{2}\pi/{5},0,\pi,-{\pi}/{5}\right)$. To validate the numerical minimization, we calculated the minimal sum of the entropies on a large number of states chosen randomly [16, 17] with Haar measure (Fig. 2), confirming that no state beats either bound (1), (2). We also checked that the sum of any two entropies never beats the Maassen-Uffink bound [8] of $\log_{2}5$ (Fig. 3) and that the sum of $four$ entropies has the unique lower bound found in Ref. [9] for every choice of the MUBs, confirming that the effect presented here requires exactly three MUBs.
@@
Figure 2: Monte-Carlo evaluation of the sums of entropies in (1) and (2) on states chosen randomly using the Haar measure. The histograms (250 bins) represent the number of states whose sum of entropies of three maximally complementary observables are equal to the value in the abscissa. (a) Simulation over $10^{9}$ random states. (b) Detail of the tails of the distributions for a simulation over $10^{10}$ random states. The orange (blue) curves refer to the entropies of $A,B,C$ ($D,E,F$). The black dashed line is at $52\log_{2}$ 5 and is approached by the left tail of the orange distribution, the blue dashed line approaches the lower bound of (2). The dotted lines are the (matching) average values of the two distributions at $\sim 5.55$.
@@
Figure 3: (a) Histograms of the Monte-Carlo evaluation of the sum of the entropies of two MUBs in dimension $d=5$ (250 bins over $10^{9}$ Haar distributed random states). As expected, all histograms match (here the histograms refer to $A,B$ in orange and $C,D$ in blue). The lower bounds of the left tails (orange-blue vertical dashed lines) approach the Maassen-Uffink bound $\log_{2}5$ (black dashed line). (b) Histograms (250 bins over $10^{9}$ states) of the sum of variances of three MUBs, evaluated over $A,B,C$ (orange) with lower bound 1.67, and $D,E,F$ (blue) with bound 1.37.
@@
We inspected all the $\binom{6}{3}=20$ possible combinations of three MUBs out of 6 and found only the two bounds reported in (1), (2). Moreover, if a triplet could attain one bound, the remaining triplet would obtain the other bound. In particular, the triplets that only achieve the bound $52\log_{2}$ 5 are $ABC,$ $ABF,$ $BEF,$ $ADE,$ $BCD,$ $CEF,$ $CDF,$ $BDE,$ $ACE,$ $ADF$ ; whereas the ones that achieve $4.43223$ are $ABD,$ $ABE,$ $ACF,$ $BCE,$ $DEF,$ $CDE,$ $BDF,$ $ACD,$ $BCF,$ $AEF$.
@@
Interestingly, even though it is known that there are two inequivalent classes of MUBs also in dimension $d=4$ [14], we could not find a similar mismatch in the sum of entropies of three MUBs: for $d=4$ there are 5 MUBs connected to five observables ($\hat{A},\hat{B},\hat{C},\hat{D},\hat{E}$), and we checked all $\binom{5}{3}=10$ combinations of triplets $\hat{X},\hat{Y},\hat{Z}$ in $\{\hat{A},\hat{B},\hat{C},\hat{D},\hat{E}\}$. They all have the lower bound $H(\hat{X})+H(\hat{Y})+H(\hat{Z})\geqslant 3$ identified in Ref. [9]. However, while in the case $d=3$ the states achieving the lower bound are the same for every triplet, in $d=4$ the optimal states depend on the choice of the MUBs. We note that they share a common structure, i.e. they all have two non-null components.
@@
By studying the sum of variances in $d=4$ and $d=5$ we find the same behavior of the uncertainty relation as the one found for the EURs: In $d=4$ the optimal states attain the same minimum for every choice of the bases, while in $d=5$ we find two distinct lower bounds dividing the triplets of MUBs in the same two sets as for the EURs. For $d=4$ all the uncertainty relations for every choice of the three MUBs read
@@
with $\hat{X}\neq\hat{Y}\neq\hat{Z}\in\{\hat{A},\hat{B},\hat{C},\hat{D},\hat{E}\}$.
For $d=5$, the uncertainty relation corresponding to the EUR in Eq. (1) reads
@@
with $XYZ$ in $\{ABC$ $ABF,$ $BEF,$ $ADE,$ $BCD,$ $CEF,$ $CDF,$ $BDE,$ $ACE,$ $ADF\}$; while the one related to the triplets in Eq. (2) is
$\Delta D^{2}+\Delta E^{2}+\Delta F^{2}\gtrsim 1.37,$ (5)
which bounds also all the remaining triplets.
@@
II Experimental verification
We experimentally tested the entropic results presented in the previous section by encoding information in photonic time-frequency modes [18]. Namely, we consider a Hilbert space generated by broadband frequency bins and their superpositions encoded in coherent light pulses. In this encoding alphabet, the five-dimensional computational basis $|a_{i}\rangle$ (associated to observable $\hat{A}$) is defined as a set of five Gaussian-shaped frequency bins centered at different frequencies, and the MUBs are generated by superimposing the fundamental bins with different phases. We note that the chosen alphabet falls in the category of pulsed temporal modes as the superposition states overlap in both time and frequency.
Mathematically, we can describe the mQPG operation as a positive-operator-valued measure (POVM) $\{\pi^{\gamma}\}$, where each POVM element $\pi^{\gamma}$ describes the measurement operator of a single channel set to detect mode $\gamma$ [20, 21, 19]. Ideally $\pi^{\gamma}=\ket{\gamma}\!\bra{\gamma}$ |; however, experimental imperfections lead to systematic errors in the POVMs, necessitating the more general description $\pi^{\gamma}=\sum_{ij}m^{\gamma}_{ij}\ket{a_{i}}\bra{a_{j}}$ |, with |$\ket{a_{i}}$ and |$\ket{a_{j}}$ eigenstates of the computational basis. When measuring a pure input state $\rho^{\xi}=\ket{\xi}\!\bra{\xi}$ | from the chosen five-dimensional Hilbert space, we will obtain output $\gamma$ with probability $p^{\gamma\xi}=\mathrm{Tr}(\rho^{\xi}\pi^{\gamma})$. In the measurement process, we use an mQPG with five channels, which can be programmed to perform projections onto any arbitrary MUB in the selected Hilbert space by assigning to each channel an eigenstate of that basis. Then, for each probed input state |$\ket{\xi}$, we calculate the entropy in each measurement basis by estimating the probability $p_{i}$ for each measurement outcome $i$ as the normalized counts in the corresponding channel.
[…]
Figure 5: Sum of the entropies calculated in $d=5$ for the two MUB triplets $CDF$ (left) and $ABE$ (right) for different types of input states: eigenstates of the MUBs in the selected triplet (green squares), eigenstates of the other MUBs (blue diamonds), random states (yellow circles), and low-entropy states that violate the previous assumption of bound (1) (red triangles). The filled markers show the experimental data, whereas the hollow markers describe the predicted results based on the characterized imperfect POVMs. The dashed green line and dash-dotted red line indicate the two lower bounds (1) and (2), respectively.

### v4

@@ Abstract
Quantum uncertainty relations impose fundamental limits on the joint knowledge that can be acquired from complementary observables: perfect knowledge of a quantum state in one basis implies maximal indetermination in all other mutually unbiased bases (MUBs). Uncertainty relations derived from joint properties of the MUBs are generally assumed to be uniform, irrespective of the specific observables chosen within a set. In this work, we demonstrate instead that the uncertainty relations can depend on the choice of observables. Through both experimental observation and numerical methods, we show that selecting different sets of three MUBs in a 5-dimensional quantum system results in distinct uncertainty bounds, i.e. in varying degrees of complementarity, in terms of both entropy and variance.
@@
Different complementary properties (MUBs) can be grouped and classified in terms of equivalence classes of complex Hadamard matrices [6, 7, 8]. We say that two sets of MUBs are equivalent if one can be mapped into the other by unitary transformations, permutations and phase factor multiplications. In this context, it was pointed out in Ref. [8] that in dimension $d=5$ there are two inequivalent classes of triplets of MUBs. For larger dimensions, a rich and involved structure of inequivalent sets of MUBs is known to exist [7, 9, 10]. In Ref. [11], a notion of operational inequivalence of sets of MUBs has been introduced, such that the inequivalence is detected whenever distinct sets of MUBs feature a different noise robustness, used as an incompatibility quantifier.
The existence of inequivalent sets of MUBs was shown to have practical consequences on Quantum Random Access Codes, since it prevents to generalize the analytic expression of the optimal average success probability for protocols $2^{d}\to 1$ to protocols $n^{d}\to 1$ with $n\geq 3$ [12]. Applications have also been found in entanglement detection based on complementary observables: in Ref. [13] it was shown that the lower bound on a given correlation function, used as an entanglement witness, depends not only on the dimension and the number of measurements, but also on the choice of the specific set of MUBs.
While these contributions were mainly focused on proving and/or detecting the existence of inequivalent sets of MUBs, here we present a fundamental implication of this: their impact on the uncertainty relations (URs). The URs are the very expression of the complementary properties of quantum observables and yield a direct link between experimental observations and quantum theory. While providing experimental validation of the URs for MUBs and high-dimensional systems, we detected different minimal uncertainties for distinct triplets of MUBs in $d=5$. Then, the inequivalence proved in Ref. [8] unvealed a finer structure of quantum URs: The information that a system can possess on three of the six complementary properties depends on which of the properties are considered. We show this by calculating the lower bound of the sum of the entropies or of the variances of the measurement outcomes of three complementary observables, complemented with Monte-Carlo simulations, and prove that such a lower bound depends explicitly on the choice of complementary observables: Complementarity-based complementarity (Fig. 1). Most importantly, the validity of distinct URs for triplets in the same complete set of MUBs opens foundational questions, finds application in a plethora of quantum protocols and yields to experimental verification, which we report here.
@@
Figure 1: Alice and Bob are both interested in joint values of observables with MUBs as eigenstates. In dimension 5 there are 6 of them: Bob measures observables $\hat{A},\hat{B},\hat{C}$, Alice measures $\hat{D},\hat{E},\hat{F}$. Alice gets more information: she can divide the DEFDEF phase space in smaller uncertainty blocks than Bob’s, even though they are both looking at MUBs.
I Uncertainty relations for inequivalent triplets of MUBs
@@
Consider three maximally complementary observables $\hat{A},\hat{B},\hat{C}$ with mutually-unbiased eigenstates $\{|a_{j}\rangle\}_{j},\{|b_{j}\rangle\}_{j},\{|c_{j}\rangle\}_{j}$. One can have partial knowledge of all three of them if the state $|\psi\rangle$ is a nontrivial superposition when expressed in any of the bases $\{|a_{j}\rangle\}_{j},\{|b_{j}\rangle\}_{j},\{|c_{j}\rangle\}_{j}$. How much joint information on them can one obtain? We need a quantification of how uncertain the outcomes of measurements of all three observables are. This is precisely what we learn from URs. In the following, we concentrate on two quantifiers  [^1] We also considered the Renyi entropies of order 2 and 1/2, but the numerics gave inconclusive results. On the one hand, we consider the sum of the Shannon entropies $H(\hat{O})=-\sum_{j}p_{j}\log_{2}p_{j}$ of the Born probabilities $p_{j}=|\langle o_{j}|\psi\rangle|^{2}$ pertaining to the measurement of $\hat{O}$. The entropies are all positive quantities, so a small sum implies large joint knowledge. The sum of entropies for maximally complementary observables has always a nontrivial lower bound given by the entropic uncertainty relations (EURs) [15, 16, 17]. On the other hand, we consider the sum of the variances [18] of $\hat{A},\hat{B},\hat{C}$. Differently from the entropy, the variance depends also on the eigenvalues of the observables, not only on the probabilities. We will choose observables with eigenvalues equal to a permutation $P(j)$ of the basis index $i$, i.e. $\hat{A}=\sum_{j}P(j)|a_{j}\rangle\langle a_{j}|$, and then minimize the variance $\Delta A^{2}=\langle\hat{A}^{2}\rangle-\langle\hat{A}\rangle^{2}$ over the permutations, to avoid effects due to the arbitrariness of the eigenvalue assignments. Also the sum of variances has a nonzero  [^2] Instead, the product of variances, as in the traditional Heisenberg-Robertson uncertainty relations [32], is problematic because it has no nonzero lower bound when minimizing over states: the product is null already if only one of the two variances is nonzero. lower bound for maximally complementary observables [20, 21]. We now show that the UR in terms of entropies or variances of three different maximally complementary observables $\hat{A},\hat{B},\hat{C}$ depends on the choice of the maximally complementary observables.
@@
We start from dimension $d=5$ where there exist six maximally complementary observables $\hat{A},\hat{B},\hat{C},\hat{D},\hat{E},\hat{F}$ with eigenvectors equal to the six MUBs [8, 7] presented in the appendix. In this case, we find that
@@
The bound (1) was known [17] and was implicitly assumed to hold for every possible choice of MUB triplet, whereas the bound (2), obtained numerically (see in the appendix), was found while experimentally testing the EURs. The states that minimize the first bound are eigenstates of any of the three $\hat{A},\hat{B},\hat{C}$, e.g. $|\psi\rangle=|a_{0}\rangle$. The states that minimize the second bound, instead, have a single null component when expressed in the computational basis $|a_{j}\rangle$, e.g. a state $\sum_{j}\psi_{j}e^{i\phi_{j}}|a_{j}\rangle$ with $\{\psi_{j}\}=(0.54488,0.45067,0,0.45067,0.54488)$ and $\{\phi_{j}\}=\left(0,-{2}\pi/{5},0,\pi,-{\pi}/{5}\right)$.
To validate the numerical minimization, we calculated the minimal sum of the entropies on a large number of pure states chosen randomly [22, 23] with Haar measure (Fig. 2), confirming that no state beats either bound (1), (2). We also checked that the sum of any two entropies never beats the Maassen-Uffink bound [16] of $\log_{2}5$ (Fig. 3) and that the sum of fourfour entropies has the unique lower bound found in Ref. [17] for every choice of the MUBs, confirming that the effect presented here requires exactly three MUBs.
@@
Figure 2: Monte-Carlo evaluation of the sums of entropies in (1) and (2) on pure states chosen randomly using the Haar measure. The histograms (250 bins) represent the number of states whose sum of entropies of three maximally complementary observables are equal to the value in the abscissa. (a) Simulation over $10^{9}$ random states. (b) Detail of the tails of the distributions for a simulation over $10^{10}$ random states. The orange (blue) curves refer to the entropies of $A,B,C$ ($D,E,F$). The black dashed line is at $2\log_{2}5$ and is approached by the left tail of the orange distribution, the blue dashed line approaches the lower bound of (2). The dotted lines are the (matching) average values of the two distributions at $\sim 5.55$.
@@
Figure 3: (a) Histograms of the Monte-Carlo evaluation of the sum of the entropies of two MUBs in dimension $d=5$ (250 bins over $10^{9}$ Haar distributed random pure states). As expected, all histograms match (here the histograms refer to $A,B$ in orange and $C,D$ in blue). The lower bounds of the left tails (orange-blue vertical dashed lines) approach the Maassen-Uffink bound $\log_{2}5$ (black dashed line). (b) Histograms (250 bins over $10^{9}$ states) of the sum of variances of three MUBs, evaluated over $A,B,C$ (orange) with lower bound 1.67, and $D,E,F$ (blue) with bound 1.37.
@@
We inspected all the $\binom{6}{3}=20$ possible combinations of three MUBs out of 6 and found only the two bounds reported in Eqs. (1) and (2). Moreover, if a triplet could attain one bound, the remaining triplet would obtain the other bound. In particular, the triplets that only achieve the bound $2\log_{2}5$ are the ones in the set
$\mathcal{S}_{1}\equiv\{$ $ABC,ABF,BEF,ADE,BCD,CEF,CDF,BDE,$
$ACE,ADF\};$ (3)
whereas the ones that achieve the bound in Eq. (2) are the triplets in the set
$\mathcal{S}_{2}\equiv\{$ $ABD,ABE,ACF,BCE,DEF,CDE,BDF,ACD,$
$BCF,AEF\}.$ (4)
$\{l_{1},l_{2},l_{3}\}$ $\psi_{0}$ $\psi_{1}$; $\phi_{1}$ $\psi_{2}$; $\phi_{2}$ $\psi_{3}$; $\phi_{3}$ $\psi_{4}$; $\phi_{4}$
ABDABD 0.23 0.67 0.67 0.23 0
$4\pi/5$ $\pi/5$ $\pi$ 0
ABEABE 0.23 0 0.23 0.67 0.67
0 $\pi$ $\pi/5$ $4\pi/5$
ACDACD 0 0.23 0.67 0.67 0.23
0 $4\pi/5$ $\pi/5$ $\pi/5$
ACFACF 0.67 0.23 0.23 0.67 0
$3\pi/5$ $4\pi/5$ $\pi/5$ 0
AEFAEF 0.23 0.67 0.67 0.23 0
$2\pi/5$ $\pi/5$ $3\pi/5$ 0
BCEBCE 0.45 0.45 0.55 0 0.55
$\pi/5$ $2\pi/5$ 0 $\pi$
BCFBCF 0.45 0.45 0.55 0 0.55
$3\pi/5$ $4\pi/5$ 0 $3\pi/5$
BDFBDF 0.45 0.45 0.55 0 0.55
$\pi$ $4\pi/5$ 0 $\pi/5$
CDECDE 0.55 0.55 0.45 0 0.45
$\pi/5$ $\pi/5$ 0 $4\pi/5$
DEFDEF 0.55 0.45 0 0.45 0.55
$2\pi/5$ 0 $\pi$ $\pi/5$
Table 1: Optimal states saturating the bound in Eq. (2). This bound can be achieved only by the triplets in $\mathcal{S}_{2}$, listed here in the first column. We report one state for each triplet, having set $\phi_{0}=0$.
We report in Table 1 a list of optimal states saturating the bound in Eq. (2), one for each triplet in $\mathcal{S}_{2}$. We show in the appendix that these states are all related to each other by specific unitary transformations. Moreover, we prove there that triplets belonging to the same set feature the same entropic uncertainty.
@@ 0 $4\pi/5$ $\pi/5$ $\pi/5$
In the case $d=4$, there is a three-parameter family of triplets of MUBs [8], but just for a specific choice of the parameters a triplet can be extended to the unique complete set, considered in Ref. [17]. Triplets selected from the complete set are all equivalent. Indeed, in this case we could not find a similar mismatch in the URs for three MUBs: in the complete set, for $d=4$ there are five MUBs connected to five observables ($\hat{A},\hat{B},\hat{C},\hat{D},\hat{E}$), and we checked all $\binom{5}{3}=10$ combinations of triplets $\hat{X},\hat{Y},\hat{Z}$ in $\{\hat{A},\hat{B},\hat{C},\hat{D},\hat{E}\}$. They all have the lower bound $H(\hat{X})+H(\hat{Y})+H(\hat{Z})\geqslant 3$ identified in Ref. [17]. It is worth mentioning that, while in the case $d=3$ the states achieving the lower bound are the same for every triplet, in $d=4$ the optimal states depend on the choice of the MUBs. We note that they share a common structure, i.e. they all have two non-null components.
Interestingly, if one does not fix the parameters of the family of triplets of MUBs in $d=4$ to get a complete set, a richer scenario of inequivalent triplets depending on three parameters can be explored. Then, different URs may be obtained, with lower bounds presumably depending on the parameters. We will not explore this case here, as our analysis focuses on the complementary properties of a same set of MUBs, but we point it out as a relevant issue for further research.
@@
By studying the sum of variances in $d=4$ and $d=5$, we find URs featuring the same behavior as the one found for the EURs: In $d=4$ the optimal states attain the same minimum for every choice of the bases, while in $d=5$ there are two distinct lower bounds dividing the triplets of MUBs in the same two sets as for the EURs. For $d=4$ all the URs for every choice of the three MUBs read
@@
[…]
For $d=5$, the UR corresponding to the EUR in Eq. (1) reads
@@
with $XYZ\in\mathcal{S}_{1}$ while the one related to the triplets in Eq. (2) is
$\Delta X^{2}+\Delta Y^{2}+\Delta Z^{2}\gtrsim 1.37,$ (7)
with $XYZ\in\mathcal{S}_{2}$.
Figure 4: Schematic of the experimental setup. The signal ($1545\text{\,}\mathrm{nm}$) and pump ($860\text{\,}\mathrm{nm}$) pulses are generated by a Ti:Sapphire ultrafast laser with an optical parametric oscillator (OPO) at a repetition rate of $80\text{\,}\mathrm{MHz}$. Two waveshapers generate the frequency-bin states in input from the signal pulse and the frequency-bin basis for the measurement from the pump pulse, respectively. In the mQPG waveguide, the signal modes are up-converted into a distinct output frequency based on their overlap with each pump mode. The output beam ($552\text{\,}\mathrm{n}\mathrm{m}$) is separated from the unconverted signal and pump beams by a shortpass filter (SP) and then detected by a commercial CCD spectrograph (Andor Shamrock 500i).
II Experimental verification
We experimentally tested the entropic results presented in Eqs. (1) and (2) by encoding information in photonic time-frequency modes [24]. Namely, we consider a Hilbert space generated by broadband frequency bins and their superpositions encoded in coherent light pulses. In this encoding alphabet, the five-dimensional computational basis $|a_{j}\rangle$ (associated to observable $\hat{A}$) is defined as a set of five Gaussian-shaped frequency bins centered at different frequencies, and the MUBs are generated by superimposing the fundamental bins with different phases. We note that the chosen alphabet falls in the category of pulsed temporal modes as the superposition states overlap in both time and frequency.
@@ II Experimental verification
A schematic of the experimental setup is shown in Fig. 4. The signal and pump pulses, centered at $1545\text{\,}\mathrm{nm}$ and $860\text{\,}\mathrm{nm}$ respectively, are generated by a combination of a Ti:Sapphire ultrafast laser with an optical parametric oscillator (OPO) at a repetition rate of $80\text{\,}\mathrm{MHz}$. The signal pulse is shaped by a commercial waveshaper to generate the frequency-bin states in input, whereas the pump pulse is shaped by a in-house-built 4f-waveshaper to generate the frequency-bin basis for the measurement. Both beams are coupled into the mQPG waveguide, where the signal modes are up-converted into a different output channel (corresponding to a distinct output frequency) based on their overlap with each pump mode, i.e., with each eigenstate of the chosen measurement basis. The output beam, centered at around $552\text{\,}\mathrm{n}\mathrm{m}$, is separated from the unconverted signal and pump beams by a shortpass filter and then detected by a commercial CCD spectrograph (Andor Shamrock 500i). The number of counts detected at each output frequency indicates the number of photons measured in the corresponding mode.
Figure 5: Sum of the entropies calculated in $d=5$ for the two MUB triplets CDFCDF (left) and ABEABE (right) for different types of input states: eigenstates of the MUBs in the selected triplet (green squares), eigenstates of the other MUBs (blue diamonds), random states (yellow circles), and low-entropy states that violate the previous assumption of bound (1) (red triangles). The filled markers show the experimental data, whereas the hollow markers describe the predicted results based on the characterized imperfect POVMs. The dashed green line and dash-dotted red line indicate the two lower bounds (1) and (2), respectively.